# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v3

@@ 4.5 Code availability
The code developed in this study is written in the Python programming language using Keras/TensorFlow (Python) libraries. For training and testing of deep learning networks, we have used an NVIDIA cluster with $4$ GPUs and $64$ GB RAM memory. The code is publicly available in https://github.com/ece7048/3Dsulci.
@@ 5.1 Classifiers performance for presence or absence of paracingulate sulcus
This discrepancy in performance between the left and right hemisphere was to be expected for two reasons. First, the paracingulate sulcus is more prominent in the left than in the right hemisphere Paus et al. (1996); Yücel et al. (2001), including in psychopathological situations such as schizophrenia Garrison et al. (2015), implying that the pattern recognition problem is better defined on the left hemisphere. Moreover, the left paracingulate sulcus shows more associations with regional cortical thickness and sulcal depth than the right one Fornito et al. (2006), implying more covariability of anatomical features contained in either of our input modalities with the presence of the paracingulate sulcus in the left hemisphere than in the right one.
@@ 5.2 Global explanations and their PCA component results
In this study, we extracted and evaluated the explainable results from both networks to avoid biased observations and to investigate whether there is a clear cause-and-effect relationship between the quality of explanation and prediction performance. The underlying assumption is that better explanations could lead to better prediction performance. We presented the results of the simple-MHL network in the main manuscript, as it showed slightly higher performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Extended Data Fig. 2 and 3).
@@
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey/white surface inputs show significant focus on the medial surface of the frontal lobe, thalamus and cingulate gyrus for detecting the presence of the paracingulate sulcus (PCS). Conversely, the absence of paracingulate sulcus (noPCS) is primarily detected by focusing on the thalamus and the anterior frontal lobe. In the left hemisphere significant focusing occurs on the thalamus following by the anterior front lobe. In the right hemisphere the region of significant focus is the thalamus and the anterior front lobe, following by the cingulate gyrus in some cases (PCA-Shape) (refer to Fig. 4).
@@
Concerning the sulcal skeleton inputs on the left and right sides, Fig. 5.a,b display the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network. The neural network distinctly emphasize key regions in both hemispheres. In the left hemisphere, there’s notable focus on superior temporal sulcus, insula, and inferior central and precentral sulci for the PCS class (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on superior temporal sulcus, collateral fissure, calcarine fissure, parieto-occipital fissure, and sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere highlights specific areas, including superior temporal sulcus, insula, and inferior central and precentral sulci (Fig. 5a.). In the right hemisphere, the focus shifts to superior temporal sulcus, inferior central sulcus, internal parietal sulcus, and anterior cingulate sulcus (Fig. 5b.).
@@
To validate the pattern indicating the presence or absence of the paracingulate sulcus, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Extended DATA Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey/white surface inputs, though the primary regions of focus remained consistent. Based on preliminary analysis (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam), we found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to when we used the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
@@
Figure 4: Simple-3D-MHL results on the left and right hemisphere of white-grey mater brain inputs. a,b, show the explainability results for the PCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam results (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views.
@@
Figure 5: Simple-3D-MHL results on the left and right hemisphere of sulcal skeleton brain inputs. a-b, show the explainability results for the noPCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views.
5.3 Global explanations from the 3D-Framework and the pattern learning results using grey/white surface inputs
For the simple-3D-MHL on grey/white surface inputs (Fig. 6a.), the right hemisphere shows focus on the medial view, in the thalamus and cingulate gyrus both in the PCS and - to a lesser extent - in the noPCS conditions. The left hemisphere, shows different focuses based on the PCS and noPCS conditions. The PCS decision mostly relies on the inferior temporal lobe and anterior inferior frontal lobe in the lateral view, and on the thalamus, both the anterior and posterior parts of the cingulate gyrus, the anterior frontal lobe, and the temporal lobe in the medial view. The noPCS condition, on the other hand, shows predominant focus on the inferior prefrontal lobe, and the temporal lobe in the lateral view, and on the thalamus and cingulate gyrus in the medial view.
For the simple-3D-CNN on grey/white surface inputs (refer to Supplementary material Extended DATA Fig. 6 a.), in the right hemisphere, PCS condition, the highlights in the lateral view are in the inferior frontal lobe and inferior temporal lobe in the PCS condition, and those in the medial view are in the thalamus, cingulate gyrus, anterior occipital lobe, and posterior temporal lobe. Still on the right hemisphere, in the noPCS condition, the focus is on the medial view in the thalamus, the cingulate gyrus, the frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focus on the same regions, laterally in the middle frontal lobe, and inferior and superior parietal lobe, and medially in the thalamus and in the cingulate, frontal and parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focuses on the lateral view of the posterior temporal lobe.
For both networks, in right hemisphere, the focus is essentially on the medial view (except for the PCS condition in the right hemisphere), with the main contributions in the thalamus and cingulate gyrus, suggesting a rather local explainability of the paracingulate sulcus’s presence. Conversely, on the left hemisphere, the focus is much more globally distributed, with strong contributions stemming from the lateral view, suggesting that the developmental mechanisms leading to the presence of the paracingulate sulcus affect the brain at a larger scale Fornito et al. (2006).
@@
Figure 6: Explainable methods scores and the extraction of the total overlapping pattern learning explanation of the left and right hemisphere and different inputs with expert’s observation. a shows the total overlapping pattern learning results for the right and left hemisphere of the brain for grey/white surface images of the simple-3D-MHL network. b shows the total overlapping pattern learning results for the right and left hemisphere of the brain for sulcal skeleton inputs of the simple-3D-MHL network. c shows the pattern learning results from the overlapping of the two deep learning networks (simple-3D-MHL, simple-3D-CNN), on the total overlapping pattern learning results of the lateral and medial view, based on experts’ observation. For the grey/white surface input we used the acronyms, Is: inferior surface, As: anterior surface, Ps: posterior surface, T: thalamus, H: hypothalamus, Fl: Frontal lobe, Ol: occipital lobe, Tl: temporal lobe, Pl: parietal lobe, Cc: corpus callosum, CG: cingulate gyrus. For the skeleton sulcal input we used the acronyms, STS: superior temporal sulcus, CF: calcarine fissure, POF: parieto-occipital fissure, SPS: sub-parietal sulcus, ICS: inferior central sulcus, CS: central sulcus, NA: none.
@@
A consistent method to overlay the outputs of the sulcal skeleton similar to the process described for the grey/white surface outputs was applied. For the simple-3D-MHL on sulcal skeleton inputs (Fig. 6b.), in the right hemisphere, the PCS and noPCS show similar trends, with lateral focuses on the superior temporal sulcus and the inferior precentral sulcus, and medial focuses on the interior parietal sulcus, the sub-parietal sulcus and both the inferior and superior rostral sulci (dimmed in the PCS condition). In the left hemisphere, the PCS and noPCS conditions show very different highlights. In the PCS condition, the main lateral focuses are in the posterior part of the inferior temporal sulcus, the occipital sulcus and the inferior frontal sulcus, and the medial focuses are in the rostral sulci (superior and inferior) and in the parietal sulci (interior, sub-parietal and temporo-occipital fissure). In the noPCS condition, the main crontributions are in the superior temporal sulcus and precentral sulcus in the lateral view, and in the sub-parietal sulcus in the medial view.
For the simple-3D-CNN on sulcal skeleton inputs (see Supplementary material Extended DATA Fig. 6 b.) both the right and the left hemispheres show similar trends in the PCS and noPCS conditions. In the right hemisphere, the lateral view is negligible compared to the medial view, where the main focus is on the callosal and parieto-occipital fissures and on the insula. In the left hemisphere, the main focus is found laterally in the superior temporal sulcus and medially in the collateral fissure and in the sub-parietal sulcus.
We thereafter identified the common regions between the two networks’ outputs and summarized them within Fig. 6c. Specifically, for the accurate detection of PCS within the skeleton sulcal hemisphere inputs, the pivotal sub-regions encompass the superior temporal sulcus, inferior precentral sulcus, calcarine fissure, parieto-occipital fissure, and sub-parietal sulcus. Conversely, when PCS is absent, the critical sub-regions within the right hemisphere skeleton sulcal hemisphere inputs encompass the superior temporal sulcus, inferior central sulcus, and sub-parietal sulcus (Fig. 6c). Transitioning to the left hemisphere, the skeleton sulcal hemisphere inputs underscore the significance of the superior temporal sulcus and sub-parietal sulcus. When PCS is not present, the important left hemisphere sulcal skeleton inputs comprise the superior temporal sulcus, inferior and central sulcus (Fig. 6c).
@@ 5.5 Ablation study of our 3D explainability framework
An ablation study in the simple-3D-MHL model was conducted to evaluate various combinations of global explanations (total-GradCam and total-SHAP) and global feature importance (total-Shape) (see Table 2). The best results were achieved by assigning the highest weight (’8’, 0.85) to total-Shape. This aligns with the idea that the feature importance of the inputs plays a crucial role in the explanations. For both hemispheres in the sulcal skeleton, a weight of ’5’, 0.5, given to total-SHAP, produced the highest faithfulness scores. In the white/grey surface, total-GradCam with the same weight yielded the best results (a weight of ’5’, 0.5). There were no significant differences observed in the complexity scores among the different combinations. The same patterns were observed for the simple-3D-CNN (Supplementary material; Table 2).
@@
To evaluate whether the global explanation from the 3D-Framework is superior to those provided by SHAP or GradCam, we scored the explanations with respect to faithfulness and complexity (see Table 1). For the simple-3D-MHL network, our proposed 3D framework outperformed total-GradCam and total-SHAP in terms of faithfulness score in the left hemisphere, with values exceeding 0.21, compared to the lower scores of less than 0.16 for total-GradCam and less than 0.11 for total-SHAP. In the right hemisphere, our proposed 3D framework again outperformed total-GradCam and total-SHAP, achieving faithfulness scores over 0.18, compared to scores of less than 0.13 for total-GradCam and less than 0.10 for total-SHAP. Regarding the complexity score, the 3D framework achieved the second-best result, with total-GradCam having the lowest score in the left hemisphere and total-SHAP in the right hemisphere.
We went one step further exploring the sulcal skeleton output of the simple-3D-MHL network as it slightly outperformed the simple-3D-CNN ((see Supplementary material Table. 2 ii.) in the classification task and deliver better global explanations faithfulness and complexity scores than the simple-3D-CNN. We applied an affine registration to the total overlapping results of sulcal skeleton inputs from each hemisphere onto a probabilistic atlas of sulci Perrot et al. (2011). Fig. 7 illustrates the distribution of the most relevant voxels for outcome decision within sulcal probabilistic areas according to different thresholds. The decisions on the right hemisphere were highly focused on specific sulci, with up to three sulci contributing to the 20% threshold, which we retained as the lower threshold (blue). Conversely, the left hemisphere predictions were based on broader considerations, with a number of sulci already contributing to the decision based on the 5% threshold, which we retained as the lower threshold for the left hemisphere (blue). For both conditions and both hemisphere, a specific focus is set on the superior temporal sulcus and its posterior branches, and a smaller but consistent contribution of the internal parietal and sub-parietal sulci. The noPCS condition on the left hemisphere additionally focuses on the precentral sulcus and the Sylvian fissure. Interestingly, no specific focus is oriented towards the internal frontal sulcus (S.F.int), the probabilistic region in which the paracingulate sulcus is located when present.
@@
Figure 7: A representation depicting the presence and absence patterns of the paracingulate sulcus on the left hemispheres of the brain’s sulcal skeleton a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Extended DATA Fig. 5., and NA: undefined.

### v4

@@ 4.5 Code availability
The code developed in this study is written in the Python programming language using pytorch, Keras, tensorflow (Python) libraries. For training and testing of deep learning networks, we have used an NVIDIA cluster with $4$ GPUs and $64$ GB RAM memory. The code is publicly available in https://github.com/ece7048/3Dsulci.
@@ 5.1 Classifiers performance for presence or absence of paracingulate sulcus
The discrepancy in performance between the left and right hemispheres was to be expected for two reasons. First, the PCS is more prominent in the left than in the right hemisphere (Paus et al. (1996); Yücel et al. (2001)), including in psychopathological conditions such as schizophrenia (Garrison et al. (2015)). Furthermore, the left PCS has a greater number of associations with regional cortical thickness and sulcal depth than the right PCS (Fornito et al. (2006)), implying more covariability of anatomical features contained in either of our input modalities with the presence of the PCS in the left hemisphere than the right.
@@ 5.2 Global explanations and their PCA component results
We extracted and evaluated the explainability results from both networks to avoid biased observations and to investigate whether there was a clear cause-and-effect relationship between the quality of explanation and prediction performance. The hypothesis was that better prediction performance could lead to better explanation. We present the results of the simple-MHL network in the main manuscript as it had slightly better performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Extended Data Fig. 2 and 3).
@@
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey-white surface inputs mainly focus on the frontal lobe (mostly inferior lateral and inferior medial), the cingulate gyrus, the temporal lobe, and occasionally the thalamus for detecting the presence or absence of the PCS. More specifically, for the detection of the presence of PCS (paracingulate sulcus; Fig. 4) in the left and right hemisphere the simple-MHL network focuses more in the frontal lobe (medial and inferior lateral), cingulate gyrus (mostly anterior), temporal lobe and sometimes thalamus. Contrary for the absence of PCS (No paracingulate sulcus; Fig. 4) in the left hemisphere the network focuses in the frontal lobe (mostly inferior medial and inferior frontal), the temporal lobe, the cingulate gyrus, occasionally the thalamus, and specifically for the PCA-GradCam, the corpus callosum. On the right hemisphere the simple-MHL network focuses in frontal lobe (medial and lateral), the temporal lobe, and the cingulate gyrus.
@@
Fig. 5.a,b displays the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network when using the sulcal skeleton inputs. The neural network shows distributed attention but still emphasizes some key regions in both hemispheres. In the left hemisphere, for the PCS class, there is focus on the superior temporal sulcus and its branches, the posterior sylvian fissure, and parts of the central and precentral sulci, with some attention on the cingulate sulcus and the medial frontal sulcus (containing the PCS) (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on the superior and inferior temporal sulci, cingulate sulcus, medial frontal sulcus (containing the PCS), and the sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere shows similar yet less specific attention than for the PCS class, with an additional focus on the sylvian fissure and insula, and the anterior cingulate sulcus (Fig. 5a.). In the right hemisphere, additional focus is shifted to the anterior cingulate sulcus, the sub-parietal sulcus, and elements of the ventricules (Fig. 5b.).
@@
To validate the pattern indicating the presence or absence of the PCS, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Extended DATA Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey-white surface inputs, although the primary regions of focus remained consistent. Based on preliminary analysis (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam), we found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
@@
Figure 4: Simple-3D-MHL results on the left and right hemisphere of grey-white surface brain inputs. a,b, show the explainability results for the PCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam results (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)).
@@
Figure 5: Simple-3D-MHL results on the left and right hemisphere of sulcal skeleton brain inputs. a-b, show the explainability results for the noPCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)).
5.3 Global explanations from the 3D-Framework and the pattern learning results using grey-white surface inputs
For the simple-3D-MHL on grey-white surface inputs (see Fig. 6a.c), in the right hemisphere the focus was on the frontal lobe, the insula, and parietal lobe in the PCS existence and, in the PCS absence (noPCS) on the temporal lobe, frontal lobe, cingulate gyrus and parietal lobe. In the left hemisphere, the PCS decision mostly relied on the cingulate gyrus, frontal lobe parietal lobe, and corpus collosum. The noPCS condition predominantly focused on the frontal lobe, and lateral temporal lobe and cingulate gyrus (see Fig. 6a.c).
For the simple-3D-CNN on grey-white surface inputs (refer to Supplementary material Extended DATA Fig. 6 a.) in the right hemisphere, highlighted the lateral inferior frontal lobe, and inferior temporal lobe in the PCS condition and the thalamus, cingulate gyrus, lateral anterior occipital lobe, and posterior temporal lobe. In the noPCS condition, the focus was on the thalamus, cingulate gyrus, the medial frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focused on the same regions: the lateral middle frontal lobe, and inferior and superior parietal lobe, and the thalamus and in the cingulate, frontal and medial parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focused on the lateral view of the posterior temporal lobe.
For both networks, in the right hemisphere the focus was primarily on the medial aspect of the brain (except for the PCS condition in the right hemisphere) with the main contributions in the frontal lobe and cingulate gyrus, suggesting a rather constrained explainability of the PCS presence. Conversely, in the left hemisphere, the focus was much more broadly distributed, with strong contributions stemming from the lateral cortex, suggesting that the developmental mechanisms leading to the presence of the PCS are related to the wider development of the brain (Fornito et al. (2006)).
@@
Figure 6: Explainable method’s scores and the extraction of the total overlapping pattern learning explanation of the left and right hemispheres and different inputs with expert’s observation. a shows the total overlapping pattern learning results for the right and left hemisphere of the brain for grey-white surface images of the simple-3D-MHL network. b shows the total overlapping pattern learning results for the right and left hemispheres of the brain for sulcal skeleton inputs of the simple-3D-MHL network. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)). c shows the pattern learning results from the overlapping of the simple-3D-MHL deep learning networks on the total overlapping pattern learning results of the lateral and medial views based on experts’ observation. For the grey-white surface input we used the acronyms, T: thalamus, H: hypothalamus, Fl: Frontal lobe, Ol: occipital lobe, Tl: temporal lobe, Pl: parietal lobe, Cc: corpus callosum, Cg: cingulate gyrus, NA: none. For the skeleton sulcal input we used the acronyms, S.T.s: superior temporal sulcus, S.T.i: inferior temporal sulcus, F.C.M.ant: anterior cingulate sulcus, S.pa.int:internal parietal sulcus, F.C.L.p: sylvian fissure, S.F.sup: superior frontal sulcus, S.F.int: internal frontal sulcus, S.F.inf: inferior frontal sulcus, S.Call: callosal sulcus.
@@
A consistent method to overlay the outputs of the sulcal skeleton, similar to the process described for the grey-white surface outputs, was applied. For the simple-3D-MHL on sulcal skeleton inputs (Fig. 6b,c), in the right hemisphere, the presence of PCS focuses on the lateral superior temporal sulcus, sylvian fissure, inferior frontal sulcus, and the anterior cingulate sulcus. The absence of PCS (noPCS; Fig. 6c) focuses on the inferior temporal sulcus, superior temporal sulcus, ventricule, callosal sulcus, and inferior frontal sulcus. In the left hemisphere, the PCS and noPCS conditions had very different highlights. In the PCS condition, the main focuses were on the lateral posterior part of the lateral superior temporal sulcus, posterior inferior temporal sulcus, sylvian fissure, and the inferior parietal sulcus. In the noPCS condition, the main contributions were in the sylvian fissure, lateral superior temporal sulcus, superior frontal sulcus, internal parietal sulcus, anterior cingulate sulcus, and ventricle.
For the simple-3D-CNN on sulcal skeleton inputs, aimed at the accurate detection of PCS within the sulcal hemisphere inputs, the pivotal sub-regions encompassed the superior temporal sulcus, inferior precentral sulcus, sylvian fissure and sub-parietal sulcus (see Supplementary Material Extended DATA Fig. 6b). Conversely, when PCS was absent, the critical sub-regions within the right hemisphere sulcal skeleton inputs encompassed the ventricule, superior temporal sulcus, internal parietal sulcus and rostral sulcus. Transitioning to the left hemisphere, the sulcal skeleton inputs underscore the significance of the superior temporal sulcus the, sylvian fissure, and the internal parietal sulcus. When PCS was not present, the important left hemisphere sulcal skeleton inputs comprised the superior temporal sulcus, ventricule, inferior precentral sulcus, internal parietal sulcus (see Supplementary Material Extended DATA Fig. 6b).
We thereafter identified the common regions between the two networks’ outputs. The overlap results of the two networks, simple-3D-CNN and simple-3D-MHL, for the presence and absence of PCS (noPCS) reveal several common regions of interest in both the right and left hemispheres. For the presence of PCS in the right hemisphere, both networks highlight the lateral superior temporal sulcus and sylvian fissure. In the left hemisphere under the PCS condition, both models emphasize the superior temporal sulcus, the sylvian fissure, and the internal parietal sulcus. For the absence of PCS (noPCS) in the left and right hemisphere, the networks overlap in highlighting the ventricule, part of superior temporal sulcus and the internal parietal sulcus.
@@ 5.5 Ablation study of our 3D explainability framework
An ablation study in the simple-3D-MHL model was conducted to evaluate various combinations of global explanations (total-GradCam and total-SHAP) and global feature importance (total-Shape) (see Table 2). The best results were achieved by assigning the highest weight (0.85) to total-Shape. This aligns with the idea that the feature importance of the inputs plays a crucial role in the explanations. For both hemispheres in the sulcal skeleton, a weight of 0.5 given to total-SHAP produced the highest faithfulness scores. For the grey-white surface inputs, total-GradCam with the same weight yielded the best results (a weight of 0.5). There were no significant differences observed in the complexity scores among the different combinations. The same patterns were observed for the simple-3D-CNN (Supplementary material; Table 2).
@@
To evaluate whether the global explanation from the 3D-Framework was superior to those provided by SHAP or GradCam, we scored the explanations with respect to faithfulness and complexity (see Table 1). For the simple-3D-MHL network, our proposed 3D framework outperformed total-GradCam and total-SHAP in terms of faithfulness score in the left hemisphere with values exceeding 0.21 compared to scores of less than 0.16 for total-GradCam, and less than 0.11 for total-SHAP. In the right hemisphere, our proposed 3D framework again outperformed total-GradCam and total-SHAP, achieving faithfulness scores over 0.18 compared to scores of less than 0.13 for total-GradCam, and less than 0.10 for total-SHAP. The 3D framework achieved the second-best result in complexity scores with total-GradCam having the lowest score in the left hemisphere and total-SHAP in the right hemisphere.
Up to this point, we have mainly explored the explanation results visually. However, we aimed to automate the process to identify the most significant subregions of interest based on the hypothesis (the classification task). To this end, we applied an affine registration to the total overlapping results of sulcal skeleton inputs from each hemisphere onto a probabilistic atlas of sulci (Perrot et al. (2011)). For this task, we explored the sulcal skeleton output of the simple-3D-MHL network as it slightly outperformed the simple-3D-CNN (see Supplementary Material Table 2 ii.) in the classification task and delivered better global explanations, faithfulness, and complexity scores than the simple-3D-CNN. Additionally, it follows patterns based on evidence from the literature (Fedeli et al. (2022); Harper et al. , Lopez-Persem et al. ).
Figure 7 presents the distribution of the most relevant voxels for outcome decisions within sulcal probabilistic areas according to different thresholds. The decisions in the right hemisphere were highly focused on specific sulci, with up to three sulci contributing to the 20% threshold, which we retained as the lower threshold (blue). Conversely, the left hemisphere predictions were based on broader considerations, with a number of sulci already contributing to the decision at the 5% threshold, which we retained as the lower threshold for the left hemisphere (blue).
For both conditions and both hemispheres, a specific focus was on the superior temporal sulcus and its posterior branches with a smaller but consistent contribution of the internal parietal and sub-parietal sulci. The noPCS condition on the left hemisphere additionally focused on the precentral sulcus and the Sylvian fissure. Interestingly, no specific focus was oriented towards the internal frontal sulcus (S.F.int), the probabilistic region in which the PCS is located when present.
@@
Figure 7: A representation depicting the presence and absence of the paracingulate sulcus on the left hemisphere sulcal skeleton input a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Extended DATA Fig. 5., and NA: undefined.